Noya drives to the city to buy software program at the computer store. Because of traffic she averaged only 15 miles per hour on the way there but averaged 35 miles per hour on the way home. If the total travel time was 2 hours how long does it take her to drive to the computer store.
Answer by nerdybill(7384)

Let x = amt of time driving to the computer store
then because "total travel time was 2 hours"
2-x = amt of time driving home
.
Using the distance formula of
d = rt
where
d is distance
r is rate or speed
t is time
.
"distance driving to the store" = "distance going home"
15x = 35(2-x)
15x = 70 - 35x
50x = 70
x = 70/50
x = 7/5
x = 1.4 hours
or 1 hour and 24 minutes